OpenStudy (anonymous):

Could there be a system of equations (one of them non-linear and the others linear) that has exactly three solutions? If so, how would you obtain one?

OpenStudy (anonymous):

Is the attached what you mean? The non-linear function is f(x) = x³ - 5x² + 10. The linear function is g(x) = x / 2 + 1. Three solutions, as required.

OpenStudy (anonymous):

No, what I mean is having a system of equations like the one below, but with three solutions: 3x-y+2z=1 (linear) x+y=-1 (linear) xy-z=0 (non-linear)

OpenStudy (amistre64):

3x-y+2z=1 x+y=-1 xy-z=0 do some substitutions .... z = xy from the bottom, sub in the first 3x-y+2xy=1 y = -1-x from the second .... sub in the first 3x-(-1-x)+2x(-1-x)=1 expand and solve the quadratic ... then test